(-2-4)(-10) Use distributive property to write the expression as an aquivalent expression.

-2x-10=-2x-10 +-4x-10
20x40=800
20+40=60 I am lost

Answer by jim_thompson5910(35256)

(-2-4)(-10) = -2 x -10 + -4 x -10 = 20 + 40 = 60